# Still More Shades of Null: An Evaluation Suite for Responsible Missing Value Imputation

### v3

@@ 3.1. Methodology for Simulating Missingness
For each dataset in our study (see Section 3.5), we designed socially-salient missingness scenarios corresponding to the three missingness mechanisms (MCAR, MAR, MNAR). Following (Joel and Doorsamy, 2022; Martínez-Plumed et al., 2019), we selected columns in which to inject null values based on the correlation coefficients between all covariates and the target variable, as well as feature importance computed using scikit-learn’s built-in functionality. We used sensitive attributes in the design of MAR missingness to simulate pre-existing bias, as described in Section 1, and injected higher rates of missingness for disadvantaged groups wherever possible (in some cases there were too few samples from disadvantaged groups), aiming to reflect realistic instances of disparate non-response caused by disparate access, distrust, or procedural injustice (Akpinar et al., 2024). Furthermore, we selected columns that may plausibly contain missingness. For example, blood pressure or cholesterol in diabetes is expected to always be observed, but family history or physical activity might be omitted or withheld.
Table 3 presents missingness scenarios for the diabetes dataset, based on the correlation coefficients and feature importances in Figure 2. Details for other datasets are available in Appendix B.2.
3.2. Missing Value Imputation (MVI) Techniques
As discussed in Section 2, there is a large number of competitive MVI techniques in the literature. For our study, we picked 8 techniques, from 5 broad categories in the taxonomy of  Emmanuel et al. (2021), namely: (1) deletion; (2) statistical techniques: median-mode and median-dummy; (3) machine learning-based techniques: miss-forest (Stekhoven and Bühlmann, 2011) and clustering (Gajawada and Toshniwal, 2012); (4) deep learning-based techniques: datawig (Biessmann et al., 2019) and auto-ml (Jäger et al., 2021); and (5) multiple imputation: boostclean (Krishnan et al., 2017). Additional details about these methods are deferred to Appendix A.1.
@@ 3.4. Shades-of-Null Architecture
The architecture of Shades-of-Null is shown in Figure 1. Our goal when developing this evaluation suite was to support the comprehensive evaluation of different MVI techniques (Section 3.2), under different missingness scenarios (Section 3.1), according to different aspects of model performance (Section 3.3). Shades-of-Null is novel in several ways. First, it is flexible and supports a variety of missingness scenarios, namely, single-mechanism, multi-mechanism and missingness shift. Second, it supports holistic evaluation of both imputation quality (including imputation fairness) and model performance — including accuracy, stability and fairness.
The central component of the architecture in Figure 1 is the benchmark controller that executes user-specified missingness scenarios by applying error injectors to input datasets. It then imputes missing values according to user-specified MVI technique(s) and pre-processes the imputed datasets using standard scaling for numerical features and one-hot encoding for categorical features. The controller then executes ML model training, including hyperparameter tuning. Finally, the evaluation module assesses imputation quality and model performance. For comprehensive model profiling, the evaluation module uses Virny (Herasymuk et al., 2024), a Python library that computes accuracy, stability and fairness metrics for a set of initialized models, for multiple sensitive attributes and their intersections. We store experimental results in MongoDB.
@@
Name Domain Size Sensitive Attributes
diabetes healthcare 905 sex
german finance 1,000 sex, age
folk-income finance 15,000 sex, race
law-school education 20,798 sex, race
bank marketing 40,004 age
heart healthcare 70,000 sex
folk-employment hiring 302,640 sex, race
Shades-of-Null incorporates two optimizations to streamline the execution of multiple experiments. Firstly, it decouples missing value imputation and model training. This enables the preservation of imputation results, including imputed datasets stored in a file system, facilitating their reuse during subsequent model training and evaluation stages. Secondly, the experimental pipeline within Shades-of-Null supports evaluation on multiple test sets simultaneously (e.g., with different rates or types of missingness), to speed up evaluation. For example, the running time of executing a pipeline with one training set and one test set is close to that of a pipeline with one train set and multiple test sets.
@@ 3.5. Datasets and Tasks
As noted in Section 1, in this work we are interested in missingness that is socially salient. With this in mind, we selected $7$ datasets from a variety of social decision-making contexts including lending, hiring, marketing, admissions, and healthcare, summarized in Table 2. For each dataset, there is a binary classification task, where being assigned a positive label corresponds to the attainment of a desirable social good (such as education, employment, or access to healthcare). We selected these datasets to have good coverage of (i) social domains and (ii) dataset sizes and other characteristics.
@@
Mechanism Missing Column Conditional Column ($I$) Pr(MS — $I$ is dis) Pr(MS — $I$ is priv)
@@
bank [^6] https://archive.ics.uci.edu/dataset/222/bank+marketing  (Moro et al., 2014) contains data from direct marketing campaigns by a Portuguese bank, between 2008 and 2013. It contains information on 40,004 potential customers, with 16 attributes (10 categorical, 6 numerical), including occupation, marital status, education, and a binary target that indicates whether the individual subscribed for a term deposit. Here, age is the sensitive attribute, with $<25$ and $>60$ as the disadvantaged group.
@@ 3.6. Model Types
We evaluate predictive performance of 6 ML models: (i) decision tree (dt__\__clf) with a tuned maximum tree depth, minimum samples at a leaf node, number of features used to decide the best split, and criteria to measure the quality of a split; (ii) logistic regression (lr__\__clf) with tuned regularization penalty, regularization strength, and optimization algorithm; (iii) gradient boosted trees (lgbm__\__clf) with tuned number of boosted trees, maximum tree depth, maximum tree leaves, and minimum number of samples in a leaf; (iv) random forest (rf__\__clf) with a tuned number of trees, maximum tree depth, minimum samples required to split a node, and minimum samples at a leaf node (v) neural network, historically called the multi-layer perceptron (mlp__\__clf) with two hidden layers, each with 100 neurons, and a tuned activation function, optimization algorithm, and learning rate; (vi) a deep table-learning method called GANDALF (Joseph and Raj, 2022) (gandalf__\__clf) with a tuned learning rate, number of layers in the feature abstraction layer, dropout rate for the feature abstraction layer, and initial percentage of features to be selected in each Gated Feature Learning Unit (GFLU) stage.
@@ 4. Single and Multi-Mechanism Missingness
To evaluate model correctness, we report results for F1, see Appendix C.1 for accuracy results. For fairness, we use binary group definitions throughout. For datasets with two sensitive attributes, we define the doubly-disadvantaged group as disadvantaged (dis) and everyone else as privileged (priv). For example: on the law-school, folk-income and folk-employment datasets, non-White women are the dis group, and White women, non-White men and White men collectively form the priv group. We report results for TPRD, and defer results for other fairness metrics to Appendix C.1. For stability, we used a bootstrap of 50 estimators, each seeing a random 80% of the training set (Efron and Tibshirani, 1994). Higher values of F1 and label stability are better, and values of TPRD close to zero are better.
@@ 4.1. Correctness of the Predictive Model
auto-ml, datawig and miss-forest are generally the best performing MVI techniques, but simpler statistical techniques (e.g., median-mode under MAR) are also competitive. This underscores the need to evaluate novel DL-based and ML-based approaches more holistically (e.g., on a variety of missingness scenarios) to ensure that they justify the additional training overhead and complexity they introduce, compared to simple statistical methods.
In line with conventional wisdom (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), we find that deletion worsens predictive performance. This effect is strongest for small datasets like diabetes, with F1 decreasing as much as 0.1 under MNAR, compared to the model trained on clean data.
@@ 4.3. Stability of the Predictive Model
Figure 5 shows label stability of models trained using different MVI techniques. In line with conventional wisdom (Domingos, 2000), stability depends most strongly on dataset characteristics (mainly size) and model type. Impute-then-classify approaches like miss-forest, auto-ml and datawig, which were the best-performing according to F1, also match the stability of the model trained on clean data under all missingness types and dataset sizes. clustering, which as was poor-performing according to F1 (correctness) is also poor-performing according to stability on small datasets (diabetes and german, with 905 and 1k samples, respectively). A notable and surprising exception is under mixed missingness on german, where clustering is competitive with other impute-then-classify approaches, despite not being competitive on MCAR, MAR or MNAR individually.
@@ 4.4. Imputation Quality and Fairness
In Figure 6, we report the imputation quality of different MVI techniques according to F1 score (for categorical columns), RMSE (for numerical columns), and KL divergence (for both numerical and categorical columns, computed over the columns with nulls only) and compare it with the F1 of the downstream model.
@@
We find that MVI techniques with vastly different imputation quality can give rise to models with comparable F1 scores, indicating that imputation correctness is not sufficiently predictive of model correctness (Zhang and Long, 2021a). For example, in Figure 6(a) on the diabetes dataset, median-dummy has an F1 score (imputation) close to 0; clustering, datawig, median-mode and deletion have an F1 score (imputation) between 0.5-0.6, and auto-ml and miss-forest have an F1 score (imputation) close to 1, but the models trained with all of these techniques have an F1 close to 1. This trend is consistent across all datasets and missingness types, and for numerical columns as well (see Figure 6(b)).
We observe similar trends for KL divergence (see Figure 6(c)), indicating that neither discrepancy nor distributional imputation quality metrics are sufficiently indicative of downstream model performance, refuting Shadbahr et al. (2023)’s claim.
@@ 5. Missingness Shift
Next, we evaluate the correctness, fairness, and stability of MVI techniques under missingness shift. We simulate missingness shift in two ways: (i) by varying the missingness mechanism between training and test (S4-9 in Table 1); and (ii) by varying the missingness rates between training and test. First, we hold the fraction of nulls in the test set constant (at 30%) and vary the fraction of nulls in the training set (10%, 30% and 50%). Then, we hold the fraction of nulls in the training set constant (at 30%) and vary the fraction of nulls in the test set ($10\%$, $20\%$, $30\%$, $40\%$ and $50\%$). Note that we have fewer settings for training missingness rates because varying the test set is less computationally demanding (as discussed in Section 3.4). We discuss results on diabetes, and defer results on other datasets, with fixed and variable training and test missingness rates, to Appendix D.
@@ 5.1. Correctness of the Predictive Model
All other MVI techniques, including boostclean, are generally robust to higher training missingness rates, and only show a 5% decrease in F1 (compared to 10% decrease with deletion), even at rates as high as 50%. A notable exception is clustering, which, surprisingly and somewhat counter-intuitively, has higher F1 at higher training missingness rates, and is actually better under MNAR than under MAR and MCAR, for all missingness rates and scenarios.
@@
Figure 45 in Appendix D.2 shows the F1 of the Random Forest model on diabetes as a function of test missingness rate. We find that F1 generally decreases with an increase in test missingness, corroborating the findings of  Shadbahr et al. (2023). A notable exception is miss-forest, which is robust to both forms of missingness shift (i.e., changing missingness mechanisms and missingness rates). As for training set missingness, F1 decreases with an increase in test missingness most steeply under MNAR (both during training and test). And, once again, clustering is an exception to this trend, instead showing invariance to test missingness rates under MNAR train (irrespective of test missingness) but not under MCAR and MAR train.
@@ 5.2. Fairness of the Predictive Model
We measured the impact of test missingness rate on fairness and found that fairness is highly sensitive to shifts in test missingness. Figure 46 in Appendix 45 shows the TPRD of the Random Forest model on diabetes as a function of test missingness rate. As before, fairness is highly sensitive to shifts in test missingness. For boostclean, datawig and auto-ml, TPRD generally increases (fairness worsens) with an increase in test error rate, although this increase is not monotonic. miss-forest is particularly robust, and although it only matches the fairness of the model trained on clean data under MCAR training, it is agnostic to increases in test missingness rate, under all missingness scenarios. For simpler methods such as deletion, clustering, median-mode and median-dummy, there is no consistent trend, even when missingness types remain the same and only missingness rates change between train and test. For example with deletion and clustering, TPRD increases with increase in test missingness rate in scenario S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but monotonically decreases with test missingness rate in scenario S2 (MAR train, MAR test).
In summary, our results corroborate the findings of Guha et al. (2024), and are a cause for concern as they indicate that the MVI techniques that perform best during development may not preserve fairness post-deployment, where shifts are likely to occur.

### v4

@@ 3.1. Methodology for Simulating Missingness
For each dataset in our study (see Section 3.5), we designed socially-salient missingness scenarios corresponding to the three missingness mechanisms (MCAR, MAR, MNAR). Following (Joel and Doorsamy, 2022; Martínez-Plumed et al., 2019), we identified features for missing value injection, denoted by $\mathcal{F}^{m}$, based on their Spearman correlation with the target variable and feature importance scores computed using scikit-learn. These selected features were chosen to reflect plausible missingness patterns. For instance, in the diabetes dataset, while features like blood pressure or cholesterol levels are expected to be consistently observed, others, such as family history or physical activity, might be omitted or withheld due to privacy concerns or reporting biases. The remaining features, denoted by $\mathcal{F}^{c}$, were considered complete, with no missing values.
The three missing mechanisms share the same set of selected features ($\mathcal{F}^{m}$), but differ in their injection strategies. For MCAR, the missing values are randomly injected on $\mathcal{F}^{m}$. In contrast, the missingness of MAR is based on sensitive attributes within $\mathcal{F}^{c}$ to simulate pre-existing bias, as described in Section 1. Specifically, higher rates of missingness were injected to disadvantaged groups wherever possible (in some cases there were too few samples from disadvantaged groups), reflecting realistic disparities caused by unequal access, distrust, or procedural injustice (Akpinar et al., 2024). Finally, for MNAR, the missingness is determined by missing values themselves, and the likelihood of missing values depends on the missing features.
Table 3 presents the selected columns ($\mathcal{F}^{m}$) and injection conditions for the diabetes dataset, based on the correlation coefficients and feature importances in Figure 2. Additional information on other datasets is available in Appendix B.2.
3.2. Missing Value Imputation (MVI) Techniques
As discussed in Section 2, many competitive MVI techniques have been proposed. We selected 15 of them, from 8 broad categories based on taxonomies presented in (Emmanuel et al., 2021; Jäger et al., 2021; Miao et al., 2022; Wang et al., 2024b), namely: (1) deletion; (2) statistical: median-mode and median-dummy; (3) machine learning-based: miss-forest (Stekhoven and Bühlmann, 2011) and clustering (Gajawada and Toshniwal, 2012); (4) discriminative deep learning-based: datawig (Biessmann et al., 2019) and auto-ml (Jäger et al., 2021); (5) generative deep learning-based: gain (Yoon et al., 2018) and hi-vae (Nazabal et al., 2020); (6) MNAR-specific: not-miwae (Ipsen et al., 2020) and mnar-pvae (Ma and Zhang, 2021); (7) multiple imputation: boostclean (Krishnan et al., 2017); and (8) other recent: nomi (Wang et al., 2024b), tdm (Zhao et al., 2023), and edit-gain (Miao et al., 2021). See Appendix A.1 for details. for details.
@@ 3.3.2. Model Performance
Broadly speaking, fairness metrics that are based on error rates are consistent with formal equality of opportunity, while those that are based on selection rates (SRD and DI) are consistent with substantive equality of opportunity (Khan et al., 2022). Which specific metric is more relevant depends on the application domain and on the stakeholder (Narayanan, 2018). SRD focuses on absolute disparities in selection rates (e.g., fixed quotas used in college admissions), while DI focuses on relative disparities (e.g., enforcing the 4/5th rule for non-discrimination in hiring in the U.S.). From the point of view of individuals being selected, TPRD is typically used when assessing fairness in opportunity allocation, such as hiring or loan approvals, where a positive outcome should be equally accessible across groups, while TNRD is most relevant in risk assessment and exclusion decisions, like medical diagnoses, where fairness in avoiding false negatives (e.g., failing to detect a disease) matters.
@@ 3.4. Shades-of-Null Architecture
The architecture of Shades-of-Null is shown in Figure 1. The core component is the benchmark controller, which executes user-specified missingness scenarios by applying error injectors to input datasets. It then imputes missing values using the selected MVI technique(s) and preprocesses the data with standard scaling for numerical features and one-hot encoding for categorical features. Next, the controller trains ML models, including hyperparameter tuning. The evaluation module then assesses imputation quality and model performance. For comprehensive profiling, it uses Virny (Herasymuk et al., 2024), a Python library that computes accuracy, stability, and fairness metrics across multiple sensitive attributes and their intersections.
@@
name domain # tuples # attrs sensitive attrs
diabetes healthcare 952 17 sex
german finance 1,000 21 sex, age
folk-income finance 15,000 10 sex, race
law-school education 20,798 11 sex, race
bank marketing 40,004 13 age
heart healthcare 70,000 11 sex
folk-employment hiring 302,640 16 sex, race
Shades-of-Null incorporates two optimizations to enhance experimental efficiency. First, it decouples missing value imputation from model training, allowing imputed datasets to be stored and reused in subsequent training and evaluation stages. Second, it supports simultaneous evaluation on multiple test sets (e.g., with varying missingness rates or types), significantly reducing running time, and so executing a pipeline with one training set and multiple test sets takes about the same time as with a single test set.
@@ 3.5. Datasets and Tasks
As noted in Section 1, we focus on socially salient missingness. With this in mind, we selected seven datasets from diverse social decision-making contexts, including lending, hiring, marketing, admissions, and healthcare, summarized in Table 2. Each dataset involves a binary classification task, where a positive label represents access to a desirable social good (e.g., education, employment, or healthcare). We chose these datasets to ensure broad coverage of (i) social domains, (ii) dataset sizes, and (iii) numerical-to-categorical column ratios. Dataset descriptions are deferred to Appendix A.3
@@
[…]
@@ 3.6. Model Types
We evaluate predictive performance of 6 ML models: (i) decision tree (dt__\__clf) with a tuned maximum tree depth, minimum samples at a leaf node, number of features used to decide the best split, and criteria to measure the quality of a split; (ii) logistic regression (lr__\__clf) with tuned regularization penalty, regularization strength, and optimization algorithm; (iii) gradient boosted trees (lgbm__\__clf) with tuned number of boosted trees, maximum tree depth, maximum tree leaves, and minimum number of samples in a leaf; (iv) random forest (rf__\__clf) with a tuned number of trees, maximum tree depth, minimum samples required to split a node, and minimum samples at a leaf node (v) neural network, historically called the multi-layer perceptron (mlp__\__clf) with two hidden layers, each with 100 neurons, and a tuned activation function, optimization algorithm, and learning rate; (vi) a deep table-learning method called GANDALF (Joseph and Raj, 2022) (gandalf__\__clf) with a tuned learning rate, number of layers in the feature abstraction layer, dropout rate for the feature abstraction layer, and initial percentage of features to be selected in each Gated Feature Learning Unit (GFLU) stage. Search grids of hyperparameters for all models are defined in our codebase.
@@ 4. Single and Multi-Mechanism Missingness
To evaluate model correctness, we report results for F1, see Appendix C.2 for accuracy results. For fairness, we use binary group definitions. For datasets with two sensitive attributes, we define the doubly-disadvantaged group as disadvantaged (dis) and everyone else as privileged (priv). For example: on the law-school, folk-income and folk-employment datasets, non-White women are the dis group, and White women, non-White men and White men are the priv group. We report results for TPRD, and defer results for other fairness metrics to Appendix C.2. For stability, we used a bootstrap of 50 estimators, each seeing a random 80% of the training set (Efron and Tibshirani, 1994). Higher values of F1 and label stability are better, and values of TPRD close to zero are better.
@@ 4.1. Correctness of the Predictive Model
auto-ml, datawig and miss-forest are generally the best performing MVI techniques, with nomi a close second, offering an optimal balance between imputation accuracy and training time (see Section 6 and Appendix C.1 for training time analysis). However, simpler statistical techniques (e.g., median-mode under MAR) are also competitive. This underscores the need to evaluate novel DL-based and ML-based methods holistically (e.g., on a variety of missingness scenarios) to ensure that they justify the additional training overhead and complexity they introduce compared to simple methods.
Interestingly, not-miwae and mnar-pvae do not demonstrate superior performance compared to other methods in our socially salient MNAR scenario. Instead, their performance aligns closely with other leading MVI approaches under MNAR conditions. This finding is further discussed in Section 7.
In line with conventional wisdom (Little and Rubin, 1987; Martínez-Plumed et al., 2019; Little et al., 2015; Joel and Doorsamy, 2022), we find that deletion worsens predictive performance. This effect is strongest for small datasets like diabetes, with F1 decreasing as much as 0.1 under MNAR, compared to the model trained on clean data. This is due to deletion discarding useful information, whereas retaining rows with nulls can still provide valuable signal for model training.
@@ 4.3. Stability of the Predictive Model
Figure 5 shows label stability of models trained using different MVI techniques. In line with conventional wisdom (Domingos, 2000), stability depends most strongly on dataset characteristics (mainly size) and model type. Impute-then-classify approaches like miss-forest, auto-ml and datawig, which were the best-performing according to F1, also match the stability of the model trained on clean data under all missingness types and dataset sizes. nomi, which performs best on accuracy and training time, demonstrates comparable stability to these top techniques on all datasets. clustering, which as was poor-performing according to F1 (correctness) is also poor-performing according to stability on small datasets (diabetes and german, with 905 and 1k samples, respectively). A notable and surprising exception is under mixed missingness on german, where clustering is competitive with other impute-then-classify approaches, despite not being competitive on MCAR, MAR or MNAR individually. The reason for this is that imputation accuracy of MVI techniques impacts data uncertainty, which then directly impacts overall model uncertainty (Gal et al., 2016).
@@ 4.4. Imputation Quality and Fairness
In Figure 6, we report the imputation quality of 10 most accurate MVI techniques per category according to F1 score (for categorical columns), RMSE (for numerical columns), and KL divergence (for both numerical and categorical columns, computed over the columns with nulls only) and compare it with the F1 of the downstream model. See Appendix C for an extended comparison of training times and accuracy across all MVI techniques.
@@
We find that MVI techniques with vastly different imputation quality can give rise to models with comparable F1 scores, indicating that imputation correctness is not sufficiently predictive of model correctness (Zhang and Long, 2021a). For example, in Figure 6(a) on the diabetes dataset, median-dummy has an F1 score (imputation) close to 0; auto-ml, miss-forest and nomi have an F1 score (imputation) close to 1; and all other imputers have an F1 score (imputation) between 0.5-0.6, but the models trained with all of these techniques have an F1 close to 1. This trend is consistent across all datasets and missingness types, and for numerical columns as well (see Figure 6(b)). We observe similar trends for KL divergence (see Figure 6(c)), indicating that neither discrepancy nor distributional imputation quality metrics are sufficiently indicative of downstream model performance, refuting Shadbahr et al. (2023)’s claim.
@@
Figure 7. Imputation fairness vs. model fairness: imputation fairness (F1 difference, RMSE difference and KL divergence difference) may not be indicative of model fairness (TPRD).
@@ 5. Missingness Shift
Next, we evaluate the correctness, fairness, and stability of 10 most effective MVI techniques from various categories under missingness shift. We simulate missingness shift in two ways: (i) by varying the missingness mechanism between training and test (S4-9 in Table 1); and (ii) by varying the missingness rates between training and test. First, we hold the fraction of nulls in the test set constant (at 30%) and vary the fraction of nulls in the training set (10%, 30% and 50%). Then, we hold the fraction of nulls in the training set constant (at 30%) and vary the fraction of nulls in the test set ($10\%$, $20\%$, $30\%$, $40\%$ and $50\%$). Note that we have fewer settings for training missingness rates because varying the test set is less computationally demanding (as discussed in Section 3.4). We discuss results on diabetes, and defer results on other datasets, with fixed and variable training and test missingness rates, to Appendix D.
@@ 5.1. Correctness of the Predictive Model
All other MVI techniques, including boostclean, are generally robust to higher training missingness rates, and only show a 5% decrease in F1 (compared to 10% decrease with deletion), even at rates as high as 50%. This is because even imperfect imputation provides valuable insights for the model, making deletion a less favorable choice. A notable exception is clustering, which, surprisingly and somewhat counter-intuitively, has higher F1 at higher training missingness rates, and is actually better under MNAR than under MAR and MCAR, for all missingness rates and scenarios.
@@
Figure 49 in Appendix D.2 shows the F1 of the Random Forest model on diabetes as a function of test missingness rate. We find that F1 generally decreases with an increase in test missingness, corroborating the findings of  Shadbahr et al. (2023) and Miao et al. (2022). A notable exception is miss-forest, which is robust to both forms of missingness shift such as changing missingness rates (as observed in (Miao et al., 2022)) and missingness mechanisms. As for training set missingness, F1 decreases with an increase in test missingness most steeply under MNAR (both during training and test), further supporting the findings of Miao et al. (2022). And, once again, clustering is an exception to this trend, instead showing invariance to test missingness rates under MNAR train (irrespective of test missingness) but not under MCAR and MAR train.
@@ 5.2. Fairness of the Predictive Model
We measured the impact of test missingness rate on fairness and found that fairness is highly sensitive to shifts in test missingness. Figure 50 in Appendix 49 shows the TPRD of the Random Forest model on diabetes as a function of test missingness rate. As before, fairness is highly sensitive to shifts in test missingness. For boostclean, datawig, and mnar-pvae, TPRD generally increases (fairness worsens) with an increase in test error rate, although this increase is not monotonic. miss-forest, auto-ml, and nomi are particularly robust to increases in test missingness rate under all missingness scenarios. For simpler methods such as deletion, clustering, median-mode and median-dummy, there is no consistent trend, even when missingness types remain the same and only missingness rates change between train and test. For example with deletion and clustering, TPRD increases with increase in test missingness rate in scenario S1 (MCAR train, MCAR test) and S3 (MNAR train, MNAR test), but monotonically decreases with test missingness rate in scenario S2 (MAR train, MAR test).
In summary,
our results corroborate the findings of Guha et al. (2024), and are a cause for concern as they indicate that the MVI techniques that perform best during development may not preserve fairness post-deployment, where shifts are likely to occur.